OpenStudy (anonymous):

need help solving \[(-2+1\sqrt{3})(5+8\sqrt{3})\]

OpenStudy (anonymous):

Expand the brackets

OpenStudy (anonymous):

i got a final answer of \[14 +4\sqrt{3}\]

OpenStudy (anonymous):

\[-10 + -16\sqrt{3} + 5\sqrt{3} + 24\]

OpenStudy (anonymous):

\[14+11\sqrt{3}\]

OpenStudy (anonymous):

Good (:

OpenStudy (anonymous):

orry read the numbers wrong

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw. take care (: